If the initial temperature of an ideal gas at 2.250 atm is 62.00 °C, what final temperature would cause the pressure to be reduced to 1.650 atm? Number
Answer #1

Apply ideal gas law ratios

P1/T1 = P2/T2

substitute known data,

change to absolute tempearture

2.25 / (62+273) = 1.65 / (T2)

T2 = 1.65/2.25*(62+273)

T2 = 245.666K

T2 = 245.666-273 = -27.334 °C